What's the solution for:
x^4 + 11x^2 - 80

Answer by ankor@dixie-net.com(22740)

To solve this you want to write it:
x^4 + 11x^2 - 80 = 0
:
Factor and solve for x^2:
(x^2 + 16)(x^2 - 5) = 0
:
x^2 = +5
x = +Sqrt(5)
x = -Sqrt(5)
These are the only real solutions
:
Did this help?